A flask is charged with 1.500 atm of N2O4(g) and 1.00 atm of NO2(g) at 25 ∘C , and the following equilibrium is achieved:

N2O4(g)⇌2NO2(g)

After equilibrium is reached, the partial pressure of NO2 is 0.519 atm .

Calculate the value of Kc for the reaction.
Answer #1

N2O4(g)⇌2NO2(g)

initially

PN2O4 = 1.500 atm

PNO2 = 1.00 atm

at equilibrium

PNO2 = 0.519 atm

1.00 - 0.519 = 0.481 atm changed

PN2O4 = 1.500 - 2(0.481) = 2.462 atm

Kp = (PNO2)2 / (PN2O4)

Kp = (0.519)2 / ( 2.462)

Kp = 0.109

we have a relation

Kp = Kc (RT)n

n = 2 - 1 = 1

0.109 = Kc (0.0821 x 298)1

0.109 = Kc 24.46

Kc = 4.46 x 10-3

We were unable to transcribe this image

We were unable to transcribe this image